# Revealing isotropic abundant low-energy excitations in UTe2 through complex microwave surface impedance

### v1

@@ B.3 Impedance fitting
$\sqrt{\frac{1}{Q(T)}\left(\frac{-2\Delta f_{0}(T)}{f_{0,\text{tot.}}(T_{0})}+x\right)}=\sum_{i=a,b,c}v_{i}R_{i}(T).$ (13)
Once the three $v_{i}$ are determined, we can use the normalization condition on $w_{i}$ to determine $G=\left(\sum_{i}v_{i}\right)^{-1}$, which allows us to convert back to the original parameters. Each measured temperature $T_{i}$ gives an instance of Eq. (13) with the same four parameters $x$, $v_{a}$, $v_{b}$, and $v_{c}$ to be determined. Since the number of measured temperatures in the normal state $N_{T,N}\sim 200$, this system is overdetermined. In matrix form, this system becomes $\bm{M}\vec{v}=\vec{F}(x)$, where $M_{ij}=R_{j}(T_{i})$ and $F_{i}(x)=\sqrt{\frac{1}{Q(T_{i})}\left(\frac{-2\Delta f_{0}(T_{i})}{f_{0,\text{tot.}}(T_{0})}+x\right)}$. We make a grid of $x$ values within the constraints discussed in §B.2 and uniquely determine $v_{i}$ which best approximates Eq. (13) in the least-squares sense for each individual value of $x$. This can result in one or more of $v_{i}$ having a negative value, which is not allowed since they are proportional to current weighting factors $w_{i}$. In such a case, we must consider induced-current weights due to a subset of the resistivity tensor components. For the case of only one resistivity tensor component, the corresponding $w_{i}=1$, so a valid solution is guaranteed. For each value of $x$, we consider the results of all seven possible combinations of one, two, and three allowed resistivity tensor components. We then choose the valid solution which results in the minimum least-squares difference between the left- and right-hand sides of Eq. (13). We finally minimize the least-squares cost function with respect to $x$ to determine the full solution. The cost function and the parameters of the fit are still continuous with respect to $x$ even when the number of allowed resistivity tensor components changes, though their derivatives are discontinuous. See the insets of Fig. (5) for an example of this optimization.
@@ C.2 Assumption-Free Determination of $\omega\tau$ in the Normal State
Here we outline a simple analysis of the normal state raw data ($f(T)$, $Q(T)$ for $T>T_{c}$) to determine $\omega\tau(T)$ using only the assumption that an isotropic single-Drude model describes the normal state electrodynamics. This analysis does not assume anything about the geometric factor $G$, the zero-temperature reactance $X_{0}$, the weighting functions $w_{i}$, etc. Reference [37] notes that a plot of raw data in the form of
@@
where $f_{BW}(T)$ is the 3-dB bandwidth ($Q=f_{0}/f_{BW}$) and $f_{0}$ is the resonant frequency, both as a function of temperature. Reference [37] shows that $\Xi(T)$ has a unique dependence on $\omega\tau(T)$ depending only on the frequency-dependent dimensionless nonlocality parameter $\beta=\frac{v_{F}}{\omega\lambda_{L}}=\frac{\ell_{mfp}/\lambda_{L}}{\omega\tau}$, where $v_{F}$ is the Fermi velocity, $\lambda_{L}$ is the London penetration depth, and $\ell_{mfp}$ is the normal-state carrier mean free path. In other words, once $\beta$ is known for a given mode, we can then convert the raw complex frequency shift data into an estimate of $\omega\tau(T)$ in the normal state. In addition, if $\beta<\beta_{c}=(1+\frac{1}{(\omega\tau)^{2}})^{3/2}$, then the sample can be considered to be in the local limit [37]. For an isotropic sample in the local limit, $\Xi(T)=X_{s}(T)/R_{s}(T)$, and $\omega\tau(T)=(\Xi^{2}(T)-1)/(2\Xi(T))$. In Fig. 16, we compare $\Xi(T)$ calculated from raw data using Eq. 14 with the ratio of $X_{s}(T)$ and $R_{s}(T)$ after the impedance fitting.

### v2

@@ B.3 Impedance fitting
$\sqrt{\frac{1}{Q_{{\color[rgb]{0,0,0}\text{sample}}}(T)}\left(\frac{-2\Delta f_{0{\color[rgb]{0,0,0},\text{sample}}}(T)}{f_{0,\text{tot.}}(T_{0})}+x\right)}=\sum_{i=a,b,c}v_{i}R_{i}(T).$ (13)
Once the three $v_{i}$ are determined, we can use the normalization condition on $w_{i}$ to determine $G=\left(\sum_{i}v_{i}\right)^{-1}$, which allows us to convert back to the original parameters. Each measured temperature $T_{i}$ gives an instance of Eq. (13) with the same four parameters $x$, $v_{a}$, $v_{b}$, and $v_{c}$ to be determined. Since the number of measured temperatures in the normal state $N_{T,N}\sim 200$, this system is overdetermined. In matrix form, this system becomes $\bm{M}\vec{v}=\vec{F}(x)$, where $M_{ij}=R_{j}(T_{i})$ and $F_{i}(x)=\sqrt{\frac{1}{Q_{{\color[rgb]{0,0,0}\text{sample}}}(T_{i})}\left(\frac{-2\Delta f_{0{\color[rgb]{0,0,0},\text{sample}}}(T_{i})}{f_{0,\text{tot.}}(T_{0})}+x\right)}$. We make a grid of $x$ values within the constraints discussed in §B.2 and uniquely determine $v_{i}$ which best approximates Eq. (13) in the least-squares sense for each individual value of $x$. This can result in one or more of $v_{i}$ having a negative value, which is not allowed since they are proportional to current weighting factors $w_{i}$. In such a case, we must consider induced-current weights due to a subset of the resistivity tensor components. For the case of only one resistivity tensor component, the corresponding $w_{i}=1$, so a valid solution is guaranteed. For each value of $x$, we consider the results of all seven possible combinations of one, two, and three allowed resistivity tensor components. We then choose the valid solution which results in the minimum least-squares difference between the left- and right-hand sides of Eq. (13). We finally minimize the least-squares cost function with respect to $x$ to determine the full solution. The cost function and the parameters of the fit are still continuous with respect to $x$ even when the number of allowed resistivity tensor components changes, though their derivatives are discontinuous. See the insets of Fig. (5) for an example of this optimization.
@@ C.2 Assumption-Free Determination of $\omega\tau$ in the Normal State
Here we outline a simple analysis of the normal state raw data (${\color[rgb]{0,0,0}\Delta}f_{{\color[rgb]{0,0,0}0,\text{sample}}}(T)$, $Q_{{\color[rgb]{0,0,0}\text{sample}}}(T)$ for $T>T_{c}$) to determine $\omega\tau(T)$ using only the assumption that an isotropic single-Drude model describes the normal state electrodynamics. This analysis does not assume anything about the geometric factor $G$, the zero-temperature reactance $X_{0}$, the weighting functions $w_{i}$, etc. Reference [37] notes that a plot of raw data in the form of
@@
where $f_{BW}(T)$ is the 3-dB bandwidth ($Q_{{\color[rgb]{0,0,0}\text{sample}}}=f_{0{\color[rgb]{0,0,0},\text{tot.}}}/f_{BW}$) of the sample and ${\color[rgb]{0,0,0}\Delta}f_{0{\color[rgb]{0,0,0},\text{sample}}}$ is the change in the resonant frequency of the sample, both as function s of temperature. Reference [37] shows that $\Xi(T)$ has a unique dependence on $\omega\tau(T)$ depending only on the frequency-dependent dimensionless nonlocality parameter $\beta=\frac{v_{F}}{\omega\lambda_{L}}=\frac{\ell_{mfp}/\lambda_{L}}{\omega\tau}$, where $v_{F}$ is the Fermi velocity, $\lambda_{L}$ is the London penetration depth, and $\ell_{mfp}$ is the normal-state carrier mean free path. In other words, once $\beta$ is known for a given mode, we can then convert the raw complex frequency shift data into an estimate of $\omega\tau(T)$ in the normal state. In addition, if $\beta<\beta_{c}=(1+\frac{1}{(\omega\tau)^{2}})^{3/2}$, then the sample can be considered to be in the local limit [37]. For an isotropic sample in the local limit, $\Xi(T)=X_{s}(T)/R_{s}(T)$, and $\omega\tau(T)=(\Xi^{2}(T)-1)/(2\Xi(T))$. In Fig. 16, we compare $\Xi(T)$ calculated from raw data using Eq. 14 with the ratio of $X_{s}(T)$ and $R_{s}(T)$ after the impedance fitting.
@@ D.1.1 Power-law penetration depth fits
Additionally, we find that, overall, there does not appear to be any correlation of $\alpha$ with $Q_{\text{sample}}$; however, the large outlier for $\alpha$ corresponds to a relatively low $Q_{\text{sample}}$, while the largest $Q_{\text{sample}}$ mode has the smallest $\alpha$. Furthermore, no apparent systematic dependence (either overall or of the outlier) of $\alpha$ on $Q_{r}$ was observed.
@@ D.2 Microwave Surface Resistance Power-Law Temperature Dependence
Similarly to $\alpha$ in §D.1.1, $\alpha_{R}$ shows no apparent overall correlation with $Q_{\text{sample}}$. The two large outliers for $\alpha_{R}$ correspond to the largest values of $Q_{\text{sample}}$. Furthermore, no apparent systematic dependence (either overall or of the outlier) of $\alpha_{R}$ on $Q_{r}$ was observed.